A grocer wants to mix peanuts and cashews to produce 20 lb bag of mixed nuts worth $6.20/lb. How many pounds of each kind of nut should she use if peanuts cost $4.80/lb and cashews cost $8/lb?
Answer by checkley71(8403)

8*X+4.8(20-X)=20*6.2
8X+96-4.8X=124
3.2X=124-96
3.2X=28
X=28/3.2
X=8.75 POUNDS OF THE $8 MIX
20-8.75=11.25 POUNDS OPF THE $4.80 MIX
PROOF
8*8.75+4.8*11.25=124
70+54=124
124=124